Incomplete handbook of all things useful

Sélène Corbineau 6 février 2026

Table des matières

1 Mechanics 3 1.1 Metric machine bolts . . . . . . . . . . . . . . . . . . . . 3

2 Electronics 4 2.1 Labelling . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 2.2 Common components . . . . . . . . . . . . . . . . . . . . 4 2.3 Topologies . . . . . . . . . . . . . . . . . . . . . . . . . . 5 2.4 Noise models . . . . . . . . . . . . . . . . . . . . . . . . . 5

3 Telecommunications 6 3.1 Amateur radio . . . . . . . . . . . . . . . . . . . . . . . . 6 3.2 Character encodings . . . . . . . . . . . . . . . . . . . . . 6

1 Mechanics

1.1 Metric machine bolts

Threads

Thread  Pitch  Tapping dia.  Clearance dia.
Close  Medium  Loose
M1.6  0.35  1.25  1.7  1.8  2.0
M2  0.4  1.6  2.2  2.4  2.6
M2.5  0.45  2.05  2.7  2.9  3.1
M3  0.5  2.5  3.2  3.4  3.6
M4  0.7  3.3  4.3  4.5  4.8
M5  0.8  4.2  5.3  5.5  5.8
M6  1  5  6.4  6.6  7.0
M8  1.25  6.8  8.4  9.0  10.0
M10  1.5  8.5  10.5  11.0  12.0

Hex heads, nuts

Thread  s  e (min.)  k (nom.)  m (max.)  m’
M1.6  3.2  3.28  1.1  1.3
M2  4.0  4.18  1.4  1.6  0.95
M2.5  5.0  5.31  1.7  2  1.35
M3  5.5  5.88  2  2.4  1.55
M4  7  7.50  2.8  3.2  1.95
M5  8  8.63  3.5  4  2.45
M6  10  10.89  4  5  2.90
M8  13  14.20  5.3  6.5  3.70
M10  16  17.59  6.4  8  4.70

2 Electronics

2.1 Labelling

Color code

Digit 0  1  2  3  4  5  6  7  8  9  -1 -2
Tol.  1% 2%  0.5% 0.25% 0.1%  5% 10%

« Ne Mangez Rien Ou Je Vous Battrai Violemment, Grand Boa. » Gold and Silver cannot occur in the mantissa.

Marking schemes With color rings, read with outermost ring to the right — 2 digits + exponent + tolerance

— 3 digits + exponent + tolerance With digits : 2 digits + exponent. The base units are Ω, µH and pF.

Preferred numbers

2.2 Common components

Complementary BJT pairs

Part n°  V_{CEO} (V) I_{C,max} (mA)  β@I_{C}  f_{T} (MHz)
BC547C,’557  45  100  420..800 @ 10  300
2N3904,’06  40  200  100..300 @ 10  300
BC327,’337  45  500  100..600 @ 100  100

Power BJTs

Part n°  V_{CEO} (V)  I_{C,max} (A)  β@I_{C}  f_{T} (MHz)
D44H11,’45  80  10  60 @ 1  50
TIP41,’42  80  6  15..75 @ 3  3

2.3 Topologies

Differential gain of diff-pair U_{d} is the one-legged bias drop.

H_{D→D} = \frac{U_{d}}{V_{T}}

2.4 Noise models

Johnson noise Typical power density is 4k_{B}T, therefore e_{n} = \sqrt{4k_{B}T R} and i_{n} = \sqrt{k_{B}T/R}.

3 Telecommunications

3.1 Amateur radio

HF frequency allocations (region I) Q-code Purity limits

3.2 Character encodings

7-bit ASCII HD4480